OpenStudy (anonymous):

Simplify the expressions below as much as possible. Leave no negative numbers under radicals and no radicals in denominators.

OpenStudy (anonymous):

\[\frac{ i \sqrt{54} }{ \sqrt{2} }\]

OpenStudy (anonymous):

\[i \sqrt{27}\]

OpenStudy (anonymous):

\[i \sqrt{25+2}\]

OpenStudy (anonymous):

\[i \sqrt{5^{2}+2}\]

OpenStudy (anonymous):

\[5i \sqrt{2}\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

by that logic, then u can pull out 2 also

OpenStudy (anonymous):

how so

OpenStudy (anonymous):

\(5i \sqrt{2}\) \(5i \sqrt{1 + 1}\) \(5i \sqrt{1^2 + 1}\) \(5i * 1 \sqrt{ 1}\)

OpenStudy (anonymous):

which is clearly WRONG !

OpenStudy (anonymous):

what

OpenStudy (anonymous):

i think answer is 3isqrt3

OpenStudy (anonymous):

when you want to simplify : \(i \sqrt{27}\) you need to write the thing inside radical as products of its factors

OpenStudy (anonymous):

9 and 3 derp

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

factors not addition

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Yes ! products of factors not addition of random numbers lol,

OpenStudy (anonymous):

@sara12345 thanks @cutestpearl

OpenStudy (anonymous):

np ;)

OpenStudy (anonymous):

yw